2. Runners World reports that the times (in minutes) of the finishers in the New York City a. Write a sentence that explains the b. Using the empirical rule, what percent of 10-km run follow the distribution N(59, 10). meaning of N(59、10). runners take more than 69 minutes to finish? c. Using the empirical rule, what percent of runners take less than 39 minutes to finish?
Answer #1

a) N(59, 10) represents that the mean time taken by finishers in the New York City 10-km run is 59 minutes with a standard deviation of 10 minutes.

b) P(X > 69)

= P(z > (69 - 59)/10)

= P(z > 1)

= 16%

c) P(X < 39)

= P(z < (39 - 59)/10)

= P(z < -2)

= 2.5%